A 6.57 kg ball rolls across the street at a rate of 7 m/s. What is the KE of the ball before it hits the curb?
1
Expert's answer
2018-04-29T13:55:08-0400
The kinetic energy
KE=(mv^2)/2=(6.57×7^2)/2=160.97 J